can you help me answer 8-5r over 6=3, -4a-1 over -10a=3/8, and y+4 over y-1=4/3

-thanks

Answer by stanbon(57407)

8-5r over 6=3,
(8-5r)/6 = 3
(8-5r) = 18
-5r = 10
r = -2
-------------------------
-4a-1 over -10a=3/8,
(-4a-1)/10a = (3/8)
Cross-multiply:
-32a - 8 = -30a
2a = -8
a = -4
=========================
y+4 over y-1=4/3
---
(y+4)/(y-1) = 4/3
---
3y+12 = 4y-4
y = 16
-----------------
Cheers,
Stan H.